Outback Outfitters sells recreational equipment. One of the company’s products, a small camp stove, sells for $150 per unit. Variable expenses are $105 per stove, and fixed expenses associated with the stove total $225,000 per month.

3.At present, the company is selling 10,000 stoves per month. The sales manager is convinced that a 10% reduction in the selling price would result in a 25% increase in monthly sales of stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

4. Refer to the data in Required 3. How many stoves would have to be sold at the new selling price to attain a target profit of $75,000 per month?

Outback Outfitters
Contribution Income Statement Present Proposed
10,000 Stoves not attempted Stoves
Total Per unit Total Per unit
Sales not attempted not attempted not attempted not attempted
Variable expenses not attempted not attempted not attempted not attempted
Contribution margin 0 $0 0 $0
Fixed expenses not attempted not attempted
Net operating income $0 $0
Answer #1

3.

Outback Outfitters
Contribution Income Statement Present Proposed
10,000 Stoves 12,500 Stoves
Total Per unit Total Per unit
Sales 1,500,000 150 1,687,500 135
Variable expenses 1,050,000 105 1,312,500 105
Contribution margin 450,000 $45 375,000 $30
Fixed expenses 225,000 225,000
Net operating income 225,000 150,000

4.

Selling price per unit = $135

Variable cost per unit = $105

Contribution margin per unit = Selling price per unit – Variable cost per unit

= 135 - 105

= $30

Units to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin per unit

= (225,000 + 75,000)/30

= 300,000/30

= 10,000 stoves

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.